# Relatives of the Hermitian curve

### v2

@@ 4 Example II - non-diagonal case
For any $(x_{0},y_{0},z_{0})\in C_{A}(\mathbb{F}_{q})$, three elements $x_{0}^{\sqrt{q}+1}$, $y_{0}^{\sqrt{q}+1}$, and $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})$ are in $\mathbb{F}_{\sqrt{q}}$. Since $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, we have
$x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=0\textrm{\ and \ }y_{0}^{\sqrt{q}+1}=0.$
Hence $x_{0}=y_{0}=0$. ∎
Example 4.4
Next we consider the matrix $A=\begin{pmatrix}\omega&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. The corresponding curve $C_{A}$ is defined by
$\omega x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (12)
Proposition 4.5
For the curve (12),
$\begin{split}C_{A}(\mathbb{F}_{q})=\{P_{\zeta}=(\zeta,1,0)|&\zeta^{\sqrt{q}+1}=-1\}\cup\{R=(0,0,1)\}\\ &\cup\{Q_{\xi\rho}=(\xi,\rho,1)|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.\end{split}$
In particular, $N_{q}(C_{A})=q+\sqrt{q}+1$ and $m_{C_{A}}=\sqrt{q}+1$.
Proof.
On the line $\{z=0\}$ [^1] During computation of examples, frequently use this type of abbreviation., the equation is $\omega(x^{\sqrt{q}+1}+y^{\sqrt{q}+1})=0.$ Hence
$C_{A}(\mathbf{F}_{q})\cap\{z=0\}=\{P_{\zeta}|\zeta^{\sqrt{q}+1}=-1.\}$
In affine space $\mathbb{P}^{2}\setminus\{z=0\}$, we may assume that the $z$-coordinate is always $1$. If $(x_{0},y_{0},1)$ is an $\mathbb{F}_{q}$-point of $C_{A}$, it satisfies
$\omega(x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1})+y_{0}^{\sqrt{q}}+y_{0}=0,$ (13)
and vice versa. Since $x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1}$ and $y_{0}^{\sqrt{q}}+y_{0}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, (13) is equivalent to
@@
Hence
$C_{A}(\mathbb{F}_{q})\cap\{z\neq 0\}=\{R\}\cup\{Q_{\xi\rho}|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.$
It is easy to show that |$|\{P_{\xi}\}|=\sqrt{q}+1$, and |$|\{R\}\cup\{Q_{\xi\rho}\}|=1+(\sqrt{q}-1)(\sqrt{q}+1).$
Example 4.6
Let $A=\begin{pmatrix}1&0&0\\ 1&0&1\\ 0&1&0\end{pmatrix},$ Hence the curve $C_{A}$ is defined by
$x^{\sqrt{q}+1}+xy^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$
Proposition 4.7
For the curve $C_{A}$ in Example 4.6, we have
$N_{q}(C_{A})=q+1,\text{\ and so\ }m_{C_{A}}=\sqrt{q}.$
Proof.
First consider the case $x=0$. Then the equation becomes $y^{\sqrt{q}}z+yz^{\sqrt{q}}=0$, which has obviously $\sqrt{q}+1$ homogeneous solutions that are $\mathbb{F}_{q}$points.
Next suppose that $x=1$. Then the equation becomes
$01+y^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=$ 0 . (14)
Let $(y_{0},z_{0})$ be any $\mathbb{F}_{q}$-solution of (14). Since $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})\in\mathbb{F}_{\sqrt{q}}$, the remaining term $1+y_{0}^{\sqrt{q}}$ is also an element of $\mathbb{F}_{\sqrt{q}}$, and so is $y_{0}^{\sqrt{q}}$. Hence $y_{0}=y_{0}^{q}=(y_{0}^{\sqrt{q}})^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}}$. Therefore the point $(y_{0},z_{0})$ satisfies
$01+y_{0}+(z_{0}+z_{0}^{\sqrt{q}})y_{0}=$ 0 (15)
from (14). Put $z_{0}+z_{0}^{\sqrt{q}}=t_{0}$, which is an elements of $\mathbb{F}_{\sqrt{q}}$. If $t_{0}\neq-1$, $y_{0}=\frac{-1}{t_{0}+1}$, which is uniquely determined by $t_{0}$. For each $t_{0}\in\mathbb{F}_{\sqrt{q}}\setminus\{-1\}$, we can find $\sqrt{q}$ solutions of $z+z^{\sqrt{q}}=t_{0}$ in $\mathbb{F}_{q}$. If $t_{0}=-1$, There are no solutions of (15) on $y_{0}$. Hence $|C_{A}(\mathbb{F}_{q})\cap\{x\neq 0\}|=\sqrt{q}(\sqrt{q}-1).$ Therefore the number of $\mathbb{F}_{q}$-points of $C_{A}$ is totally
$(\sqrt{q}+1)+\sqrt{q}(\sqrt{q}-1)=q+1.$
This completes the proof. ∎
5 Example III - non-diagonal case (2)
In this section, we handle the curve
$x^{\sqrt{q}}z+xy^{\sqrt{q}}+\omega yz^{\sqrt{q}}=0,$ (16)
where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. The number of $\mathbb{F}_{q}$-points of curves in each examples explained in previous two sections does not depend on the choice of $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, however that of the curve (16) depends on the choice of $\omega$. So we denote the corresponding matrix $\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&\omega&0\end{pmatrix}$ by $A_{\omega}$.
Lemma 5.1
Let $C_{A_{\omega}}$ be a curve defined by (16).
(i)   $C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z=0\}=\{(1,0,0),(0,1,0)\}.$
${\rm(ii)}\begin{split}(C_{A_{\omega}}(\mathbb{F}_{q})\cap&\{z\neq 0\})\cap\{y^{\sqrt{q}+1}\neq 1\}\\ &=\left\{\left(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1},y_{0},1\right)\mid y_{0}\in\mathbb{F}_{q},\ y_{0}^{\sqrt{q}+1}-1\neq 0\right\}.\end{split}$
In particular,
$N_{q}(C_{A_{\omega}})\geq q-\sqrt{q}+1.$
Proof.
The assertion (i) is obvious. Let $(x_{0},y_{0},1)\in C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\}$. Then
$x_{0}^{\sqrt{q}}+x_{0}y_{0}^{\sqrt{q}}+\omega y_{0}=0.$ (17)
Taking $\sqrt{q}$-th power of this equation, we have
$x_{0}+x_{0}^{\sqrt{q}}y_{0}+\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}=0.$ (18)
[…]
$x_{0}(y_{0}^{\sqrt{q}+1}-1)-(\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2})=0.$
Hence, when $y_{0}^{\sqrt{q}+1}-1\neq 0$, $x_{0}=\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1}$. Conversely, by straightforward computation, we get
$(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})^{\sqrt{q}}+(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})y_{0}^{\sqrt{q}}+\omega y_{0}=0.$
Putting (i), (ii) together, we have $q-(\sqrt{q}+1)+2$ points over $\mathbb{F}_{q}$. ∎
For computing the number of $\mathbb{F}_{q}$-points of $(C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\})\cap\{y^{\sqrt{q}+1}=0\},$ we fix a primitive element $\omega_{0}\in\mathbb{F}_{q}^{\ast}$.
Note that $\omega_{0}^{s}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ if and only if $s\equiv 0\mod\sqrt{q}+1$.
Lemma 5.2
(1) Suppose that $q$ is odd. Then
(i) $-1=\omega_{0}^{\frac{q-1}{2}},$ and
(ii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{\frac{\sqrt{q}+1}{2}+s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}.$
(2) Suppose that $q$ is even. Then
(iii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}$.
Proof.
(i) Since $\omega_{0}$ is of order $q-1$ in the multiplicative group $\mathbb{F}_{q}^{\ast}$ and $q-1$ is even, $\omega_{0}^{\frac{q-1}{2}}=-1$.
(ii) For $b\in\mathbb{F}_{q}^{\ast}$, $\operatorname{Tr}b=0$ means $b^{\sqrt{q}-1}=-1$. Letting $b=\omega_{0}^{l}$, we have $\omega_{0}^{l(\sqrt{q}-1)}=\omega_{0}^{\frac{q-1}{2}}$. Hence $l(\sqrt{q}-1)\equiv\frac{\sqrt{q}+1}{2}(\sqrt{q}-1)\mod q-1$, which means that $l\equiv\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1$.
(iii) Similar to the case (ii). ∎
Theorem 5.3
  1. (1)
    If $(\sqrt{q}+1,3)=1$, then
    $N_{q}(C_{A_{\omega}})=q+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}.$
  2. (2)
    When $\sqrt{q}+1$ is divisible by $3$,
    1. (a)
      if $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
      $N_{q}(C_{A_{\omega}})=q+2\sqrt{q}+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}+2,\text{\ and}$
    2. (b)
      if $\omega^{\frac{\sqrt{q}+1}{3}}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
      $N_{q}(C_{A_{\omega}})=q-\sqrt{q}+1\text{\ and \ }m_{C_{A_{\omega}}}=\sqrt{q}-1.$
Proof.
We have to count the number of $\mathbb{F}_{q}$-solutions of the simultaneous equations
[…]
By changing the sign $y\mapsto-y,$ this system is equivalent to
[…]
For a given $y_{0}\in\mathbb{F}_{q}^{\ast}$ with $y_{0}^{\sqrt{q}+1}=1$, we ask whether the equation on $x$
$x^{\sqrt{q}}-y_{0}^{\sqrt{q}}x-\omega y_{0}=0$ (19)
has an $\mathbb{F}_{q}$-solution or not. Since $\operatorname{Nm}y_{0}^{\sqrt{q}}=\operatorname{Nm}y_{0}=1$, there is an element $c\in\mathbb{F}_{q}^{\ast}$ such that $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$ by Lemma 3.1. Therefore the problem above is transformed whether the equation on $cx$
$(cx)^{\sqrt{q}}-(cx)-\omega c^{\sqrt{q}}y_{0}=0$ (20)
has an $\mathbb{F}_{q}$-solution or not. From Corollary 3.2, (20) has an $\mathbb{F}_{q}$-solution if and only if $\operatorname{Tr}(\omega c^{\sqrt{q}}y_{0})=0$, and in this case all solutions of (20) are $\mathbb{F}_{q}$-solutions.
Let $i,j,k$ be integers with $11\leq i,j,k\leq q-$ 1 satisfying
$\omega=\omega_{0}^{k},\ c=\omega_{0}^{j},\text{\ and \ }y_{0}=\omega_{0}^{i}.$
Since $y_{0}^{\sqrt{q}+1}=1$, there is a unique integer $l$ with $11\leq l\leq\sqrt{q}+$ 1 so that $i\equiv l(\sqrt{q}-1)\mod q-1$. Since $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$,
$j(1-\sqrt{q})\equiv l(\sqrt{q}-1)\sqrt{q}\mod q-1.$
Hence
$j\equiv-l\sqrt{q}\mod\sqrt{q}+1.$ (21)
Since $\omega c^{\sqrt{q}}y_{0}=\omega_{0}^{k+j\sqrt{q}+l(\sqrt{q}-1)},$ $\operatorname{Tr}\omega c^{\sqrt{q}}y_{0}=0$ if and only if
$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}\end{cases}$ (22)
by Lemma 5.2. On the other hand,
$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv k-l(q-\sqrt{q}+1)\mod\sqrt{q}+1$ (23)
by (21). Together with $\sqrt{q}\equiv-1\mod\sqrt{q}+1$ and $q\equiv-\sqrt{q}\equiv 1\mod\sqrt{q}+1,$ (22) and (23) imply
$k-3l\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}.\end{cases}$ (24)
Therefore the number of $y_{0}$’s for each of which (19) has $\mathbb{F}_{q}$-solutions, in fact $\sqrt{q}$-solutions, coincides with the number of $l$’s with $11\leq l\leq\sqrt{q}+$ 1 each of which satisfies (24) for the fixed $k$.
Let us consider the map
$\begin{array}[]{rccc}\varphi:&\mathbb{Z}/(\sqrt{q}+1)&\to&\mathbb{Z}/(\sqrt{q}+1)\\ &x&\mapsto&3x\end{array}.$
If $\sqrt{q}+1$ is not divisible by $3$, $\varphi$ is bijective. Hence there exists a unique $l$ satisfying (24) for any fixed $k$. If $3$ is a factor of $\sqrt{q}+1$, then
$|\varphi^{-1}(m)|=\begin{cases}3&\text{if $m\in\operatorname{Im}\varphi$}\\ 0&\text{else}.\end{cases}$
Hence if $k\equiv 0\mod 3$, there are three $l$’s satisfying (24), and otherwise no $l$’s. The condition $k\equiv 0\mod 3$ is equivalent to the condition $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}.$ This completes the proof. ∎
Finally we pose a question: when $q\geq 16$, does a Hermitian-relative curve $C$ with $22\leq m_{C}\leq\sqrt{q}-$ 2 exist?

### v3

@@ 3.1 Some equations
Let $\lambda=\eta^{\prime}\eta^{-1}$, and suppose that $\lambda\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Since $\operatorname{Nm}$ is surjective, there is an element $\alpha\in\mathbb{F}_{q}^{\ast}$ such that $\alpha^{\sqrt{q}+1}=\lambda$. So
$\operatorname{diag}[1,1,\alpha]^{\ast}\operatorname{diag}[1,1,\eta]\operatorname{diag}[1,1,\alpha]=\operatorname{diag}[1,1,\eta^{\prime}].$
Conversely, let $T=(a_{ij})\in GL(3,\mathbb{F}_{q})$ such that
$T^{\ast}\operatorname{diag}[1,1,\eta]T=\rho\operatorname{diag}[1,1,\eta^{\prime}],$
where we are regarding matrices as in $GL(3,\mathbb{F}_{q})$ and $\rho\in\mathbb{F}_{q}^{\ast}$. For both curves $C_{\operatorname{diag}[1,1,\eta]}$ and $C_{\operatorname{diag}[1,1,\eta^{\prime}]}$, their sets of $\mathbb{F}_{q}$-points consist of $\sqrt{q}+1$ points on the line $\{z=0\}$ by Lemma 3.11. Hence the line $\{z=0\}$ is stable by the linear transformation of $\mathbb{P}^{2}$ corresponding $T$, that is
$\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33}\end{pmatrix}\begin{pmatrix}\ast\\ \ast\\ 0\end{pmatrix}=\begin{pmatrix}\ast\\ \ast\\ 0\end{pmatrix}.$
Hence $a_{31}=a_{32}=0$, and
$\begin{pmatrix}a_{11}^{\sqrt{q}}&a_{21}^{\sqrt{q}}&0\\ a_{12}^{\sqrt{q}}&a_{22}^{\sqrt{q}}&0\\ a_{13}^{\sqrt{q}}&a_{23}^{\sqrt{q}}&a_{33}^{\sqrt{q}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\eta\end{pmatrix}\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ 0&0&a_{33}\end{pmatrix}=\rho\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\eta^{\prime}\end{pmatrix}.$
Comparing $(1,1)$-component of the both sides, we have $a_{11}^{\sqrt{q}+1}+a_{21}^{\sqrt{q}+1}=\rho$, which guarantees that $\rho$ is an element of $\mathbb{F}_{\sqrt{q}}^{\ast}$. Taking determinant of the both sides, we have
$\det T^{(\sqrt{q})}\cdot\eta\cdot\det T=\rho^{3}\eta^{\prime}.$
Hence $\eta^{\prime}\eta^{-1}=(\det T)^{\sqrt{q}+1}\rho^{-3}\in\mathbb{F}_{\sqrt{q}}^{\ast}$
Proof of Theorem 3.1.
(a) Suppose that $\omega=1$. Then by Lemmas 3.8, 3.9, the curve $C_{\omega}$ has $\sqrt{q}^{3}+1$ rational points. Hence it is Hermitian if $q\neq 4$ [1], and all rational points on a Hermitian curve are inflexions. When $q=4$, a non-Hermitian curve with $9$ rational points can’t have inflexions [3, Remark 3.10].
(b) Suppose that $\operatorname{Nm}\omega=1$ and $\omega\neq 1$. Then by Lemmas 3.8, 3.9, the curve $C_{\omega}$ has $\sqrt{q}+1$ rational points, and those points lie on the line $\{z=0\}$. Let $R$ be one of those points. If $R$ is not an inflexion, then $C_{\omega}.T_{R}(C_{\omega})=\sqrt{q}R+R^{\prime}$ with $R^{\prime}\neq R$. Since $R^{\prime}$ is $\mathbb{F}_{q}$-point too, $R^{\prime}$ lies on the line $\{z=0\}$. Hence $T_{R}(C_{\omega})$ coincides with the line $\{z=0\}$, which is impossible. Therefore all the $\sqrt{q}+1$ rational points are inflexions.
Next we show that $C_{\omega}$ with $\operatorname{Nm}\omega=1$ and $\omega\neq 1$ is projectively equivalent to
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ (18)
for a certain $\eta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}.$ Since $\operatorname{Nm}\omega=1$, there is an element $\beta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$ such that $\beta^{1-\sqrt{q}}=\omega$. Using this $\beta$, (5) can be rewritten as
$(\beta x)^{\sqrt{q}}y+(\beta x)y^{\sqrt{q}}+\beta^{\sqrt{q}}z^{\sqrt{q}+1}=0.$ (19)
Replacing $\beta x$ with the new $x$ and applying Lemma 3.10, we get the equation
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\beta^{\sqrt{q}}z^{\sqrt{q}+1}=0.$
Replace $\beta^{\sqrt{q}}$ by $\eta$, then we have (18). Conversely, tracing the reverse process, we know that the curve defined by (18) is projectively equivalent to $C_{\eta^{\sqrt{q}-1}}$.
The last statement in (b) is just Lemma 3.12.
(c) Suppose that $\operatorname{Nm}\omega\neq 1$. From Lemmas 3.8, 3.9, $C_{\omega}(\mathbb{F}_{q})$ consists of $q+1$ rational points. Note that $P=(1,0,0)$ and $Q=(0,1,0)$ are inflexions, and they are only $\mathbb{F}_{q}$-points lie on the line $\{z=0\}$ by Lemma 3.5.
We will show that any point $R\in C_{\omega}(\mathbb{F}_{q})\setminus\{P,Q\}$ is not an inflexion. Since $R$ does not lie on the line $\{z=0\}$, we may assume that $R=(x_{0},y_{0},1)$, where $(x_{0},y_{0})$ satisfies the affine equation (15), in particular $x_{0}\neq 0$. We want compute the divisor $C_{\omega}.T_{R}(C_{\omega})$. Let $f(x,y)=x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+1$. Since $f_{x}=\omega y^{\sqrt{q}}$ and $f_{y}=x^{\sqrt{q}}$, the tangent line at $R$ is given by
$\omega y_{0}^{\sqrt{q}}x+x_{0}^{\sqrt{q}}y+1=0,$
or, alternatively
$y=-\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}x-(\frac{1}{x_{0}})^{\sqrt{q}}$ (20)
because $x_{0}\neq 0$. Take $\sqrt{q}$-th power of (20), then
$y^{\sqrt{q}}=-\omega^{\sqrt{q}}(\frac{y_{0}}{x_{0}})x^{\sqrt{q}}-(\frac{1}{x_{0}}).$ (21)
Substitute (20) and (21) into (15), then
$-\left\{\left(\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}+\omega^{\sqrt{q}+1}\frac{y_{0}}{x_{0}}\right)x^{\sqrt{q}+1}+(\frac{1}{x_{0}})^{\sqrt{q}}x^{\sqrt{q}}+\frac{\omega}{x_{0}}x-1\right\}=0.$ (22)
On the other hand, since
$0=\omega\left(\frac{f(x_{0},y_{0})}{x_{0}^{\sqrt{q}+1}}\right)^{\sqrt{q}}=\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}+\omega^{\sqrt{q}+1}\frac{y_{0}}{x_{0}}+\frac{\omega}{x_{0}^{\sqrt{q}+1}},$
the coefficient of $x^{\sqrt{q}+1}$ in (22) is $\frac{\omega}{x_{0}^{\sqrt{q}+1}}$. Hence (22) can be written as
$\frac{\omega}{x_{0}^{\sqrt{q}+1}}x^{\sqrt{q}+1}-\frac{1}{x_{0}^{\sqrt{q}}}x^{\sqrt{q}}-\frac{\omega}{x_{0}}x+1$ (23)
$=$ $\frac{\omega}{x_{0}^{\sqrt{q}+1}}\left(x-x_{0}\right)^{\sqrt{q}}\left(x-\frac{x_{0}}{\omega}\right)=0.$ (24)
Since $\omega\neq 1$, $x_{0}\neq\frac{x_{0}}{\omega}$, which means that $R$ is not an inflexion. Hence only $P$ and $Q$ are inflexions of $C_{\omega}$.
Lastly we classify the curves $C_{\omega}$ with $\omega\neq 1$ up to projective equivalence. For $\omega,\omega^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\operatorname{Ker}(\operatorname{Nm})$, let
$A_{\omega}=\begin{pmatrix}0&1&0\\ \omega&0&0\\ 0&0&1\end{pmatrix}\text{ and }A_{\omega^{\prime}}=\begin{pmatrix}0&1&0\\ \omega^{\prime}&0&0\\ 0&0&1\end{pmatrix}.$
Then two curves $C_{\omega}$ and $C_{\omega^{\prime}}$ are projectively equivalent to each other if and only if there is a matrix $T=(a_{ij})\in GL(3,\mathbb{F}_{q})$ such that
$T^{\ast}A_{\omega}T=\rho A_{\omega^{\prime}}$ (25)
for some $\rho\in\mathbb{F}_{q}^{\ast}$. Since inflexions of both curves are $\{P,Q\}$ that are lie on the line $\{z=0\}$, and
$T_{P}(C_{\omega})\cap T_{Q}(C_{\omega})=T_{P}(C_{\omega^{\prime}})\cap T_{Q}(C_{\omega^{\prime}})=\{(0,0,1)\},$
the line $\{z=0\}$ and the point $(0,0,1)$ are stable by the projective transformation $T$. Hence we may choose $T$ as $\begin{pmatrix}a_{11}&a_{12}&0\\ a_{21}&a_{22}&0\\ 0&0&1\end{pmatrix}.$ Then, comparing $(3,3)$-entries of both sides of (25), we know $\rho=1$. Now we write down (25) componentwise:
$\omega a_{21}^{\sqrt{q}}a_{11}+a_{11}^{\sqrt{q}}a_{21}$ $=0$ (26)
$\omega a_{22}^{\sqrt{q}}a_{12}+a_{12}^{\sqrt{q}}a_{22}$ $=0$ (27)
$\omega a_{21}^{\sqrt{q}}a_{12}+a_{11}^{\sqrt{q}}a_{22}$ $=1$ (28)
$\omega a_{22}^{\sqrt{q}}a_{11}+a_{12}^{\sqrt{q}}a_{21}$ $=\omega^{\prime},$ (29)
where (26)–(29) come from comparison of $(1,1)$,(2,2)22(2,2)( 2 , 2 ), $(1,2)$ and $(2,1)$ components respectively. Take $\sqrt{q}$th power of (26)
$\omega^{\sqrt{q}}a_{21}a_{11}^{\sqrt{q}}+a_{11}a_{21}^{\sqrt{q}}=0,$ (30)
@@
$(\omega^{\sqrt{q}+1}-1)a_{11}^{\sqrt{q}}a_{21}=0.$
Since $\operatorname{Nm}\omega\neq 1$, we conclude that $a_{11}=0$ or $a_{21}=0$. Similarly, using (27), we have $a_{22}=0$ or $a_{12}=0$. Since $\det T\neq 0$, only possibilities of $T$ are
$T=\begin{pmatrix}a_{11}&0&0\\ 0&a_{22}&0\\ 0&0&1\end{pmatrix}\text{or}\begin{pmatrix}0&a_{12}&0\\ a_{21}&0&0\\ 0&0&1\end{pmatrix}.$
If the first case occurs, then $a_{11}^{\sqrt{q}}a_{22}=1$ by (28) and $\omega a_{22}^{\sqrt{q}}a_{11}=\omega^{\prime}$ by (29), which implies that $\omega=\omega^{\prime}$. If the second case occurs, then
$\omega a_{21}^{\sqrt{q}}a_{12}=1\text{ and }a_{12}^{\sqrt{q}}a_{21}=\omega^{\prime},$
which implies that $\omega\omega^{\prime\sqrt{q}}=1$. ∎